# Weight-Space Linear Recurrent Neural Networks

### v1

@@ 4 Related Work
The problem of sequence modeling, long dominated by Transformers Vaswani et al. (2017), is experiencing a renewed focus on recurrent architectures, particularly for their efficiency and unique modeling capabilities Gu et al. (2021); De et al. (2024). Our framework, the Weight-space Adaptive Recurrent Predictor (WARP), intersects with several active research areas.
@@
The concept of leveraging the weight-space of neural networks is not new; for instance, optimizers and hypernetworks inherently process weight-space features Andrychowicz et al. (2016); Zhou et al. (2024); Ha et al. (2016), treating them as outputs of a learning algorithm. Other works explore weight features as inputs for model analysis Unterthiner et al. (2020); Schürholt et al. (2024) or for implicitly representing data Dupont et al. (2022). WARP distinguishes itself by directly evolving the root network’s weights as its intermediate state.
@@
Linear RNNs and SSMs have re-emerged as powerful tools, largely due to their parallelizable and hardware-aware training Yang et al. (2024); Movahedi et al. (2025), with impressive performance on long sequences. Notable architectures like S4 Gu et al. (2021) and Linear Attention Katharopoulos et al. (2020) have massively catalyzed recent advancements. While WARP builds on the efficiency of linear recurrence, its core innovation lies in its unique state parameterization — rather than solely on the recurrent mechanism — which includes non-linearities for improved expressivity.
@@
The landscape of sequence modeling is rich with innovative designs. Hybrid models like Griffin De et al. (2024) merge recurrences with attention, while Movahedi et al. (2025) seek to compute dense linear RNNs from diagonal ones via fixed-point transformations. Frameworks like FACTS introduce structured memories Nanbo et al. (2025). Brain-inspired architectures Zucchet et al. (2023); Ba et al. (2016), including time-varying decoder architectures Jadhav et al. (2023), seek to learn evolving relationships between model inputs and outputs. WARP contributes to this evolving field by introducing a novel mechanism — viewing the RNN hidden state as the weights and biases of a time-varying root neural network — which results in non-linear self-decoding.
@@
A central debate revolves around the true state-tracking and memory capabilities of various recurrent architectures. While some SSMs and even Transformers face theoretical limitations in solving certain tasks Merrill et al. (2024); Jelassi et al. (2024), improvements like incorporating negative eigenvalues in linear RNNs aim to enhance state-tracking Grazzi et al. (2024). Other works explicitly include neural memory modules so that surprising events are more memorable Behrouz et al. (2024). WARP’s use of a high-dimensional weight-space for its states is a direct attempt to provide richer “infinite-dimensional” memory capacity and more expressive temporal dynamics compared to conventional compressed state representations.
@@
Effective adaptation to out-of-distribution dynamics or in continual learning settings is a significant challenge Green et al. (2024). For instance, standard Neural Ordinary Differential Equations Chen et al. (2018) struggle with distribution shifts and need retraining or fine-tuning for adaptation Kirchmeyer et al. (2022); Kassaï Koupaï et al. (2024). With its gradient-free formulation, WARP facilitates test-time generalization, a problem explored in meta-learning frameworks like Neural Context Flows Nzoyem et al. (2025a), through differentiable closed-form solvers Bertinetto et al. (2018), or in-context learning Von Oswald et al. (2023).
5 Discussion & Conclusion
5.1 Benefits
WARP demonstrates outstanding results across a multitude of data modalities, both in-distribution and out-of-distribution, as evidenced by the extensive empirical results on time series forecasting and classification we have presented (see Tables 1, 2, 2 and 3). Additional results showcasing a 93% classification accuracy on sequential MNIST, along with ablation studies and further results on synthetic datasets are provided in Appendix D.
By letting the data directly interact with the weights as in Equation 1, WARP is able to fine-tune its weights without gradients, an appealing property given recent interest in test-time adaptation Behrouz et al. (2024); Wei et al. (2022). Additionally, WARP is the latest in the growing field of Scientific Machine Learning Cuomo et al. (2022) that seamlessly integrates interpretable physical knowledge into its predictions, a feature standard RNNs have overlooked. Finally, the WARP architecture bears resemblance to synaptic plasticity in biological neural networks Caporale and Dan (2008). This neuromorphic quality enables more biologically plausible learning dynamics.
5.2 Limitations
Some design decisions that strengthen WARP equally face limitations. We outline them below as promising avenues for future work.
Scaling to larger root networks.
The size of the matrix $A$ limits scaling to huge root neural networks. Our experiments conducted on a RTX 4080 GPU with 16GB memory could only support moderate $D_{\theta}$ values, leaving open the question of how expressive WARP models can become if scaled further. Future work could also explore the opposite direction, aiming to make $A$ much smaller while preserving performance via low-rank or diagonal parametrization Gupta et al. (2022).
Limited theoretical understanding.
Although our work benefits from the theoretical underpinnings of Neural Controlled Differential Equations (CDEs) as universal approximators generalizing RNNs Kidger (2022), it remains mostly empirical. With the field of weight-space learning currently dominated by empirical work, more theoretical research is needed; for instance to understand and reduce the memory footprint of the matrix $A$ as illustrated in LABEL:fig:a_matrix, or to enforce neuron permutation invariance Zhou et al. (2023).
Extension to diverse tasks.
Present work has focused on forecasting and classification. Future work could expand the framework to cover other forms of analyses like time series imputation or control. This could be achieved potentially through different parameterizations for the root network (e.g., CNNs, RNNs, Transformers); some of which might alleviate observed difficulties with long-range dependencies (see Table 3).
5.3 Broader Impact
While their benefits are evident from Section 5.1, malicious deployment of our self-adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. Additionally, high-energy costs from high-dimensional weight-space computations could increase disparities in our field. To limit the potential for such outcomes and to improve the democratization of AI, our data and models are openly available at https://anonymous.4open.science/r/warp.
5.4 Conclusion
In this work, we introduced Weight-Space linear RNNs, a novel family of sequence models that operates directly within the weight-space of neural networks, offering a distinct paradigm from traditional recurrent architectures. We argue that the high-dimensional weight-space can be used for intermediate representations, resulting in infinite-dimensional RNN hidden states and high-capacity memory. Our comprehensive experiments demonstrate that our models exhibits superior expressivity and generalization capabilities, enabling a powerful form of gradient-free adaptation in response to sequential inputs, and showing exceptional abilities when integrating domain-specific knowledge from physical systems. Our framework draws intriguing parallels to neuromorphic learning principles, leading us a step further towards human-level artificial intelligence.
Acknowledgments and Disclosure of Funding
This work was supported by UK Research and Innovation grant EP/S022937/1: Interactive Artificial Intelligence, EPSRC program grant EP/R006768/1: Digital twins for improved dynamic design, EPSRC grant EP/X039137/1: The GW4 Isambard Tier-2 service for advanced computer architectures, and EPSRC grant EP/T022205/1: JADE: Joint Academic Data science Endeavour-2.
Appendix A Methodology
A.1 Motivation
@@
The main motivation behind WARP (Weight-space Adaptive Recurrent Prediction) is gradient-free adaptation to out-of-distribution (OoD) settings. Relative to OoD detection which has always been a central problem in machine learning spanning decades of research interest Cui and Wang [2022], Keshtmand et al. [2024], OoD adaptation is a recent but growing field rich in new and stimulating ideas Arjovsky [2020], Nzoyem et al. [2025a]. WARP mimics the dynamics of gradient descent as observed through a projection of a 4898-dimensional space into a 2-dimensional PCA space in LABEL:fig:warp_vs_overfit. This offers a promising avenue for OoD adaptation with minimal cost.
A.2 Training Algorithms
@@
WARP’s Recurrent Training Algorithm Input: Training sequences $\{\mathbf{x}^{i}_{t}\}_{t\in\{0,\ldots,T-1\}}^{i\in\{0,\ldots,N-1\}}$ Output: Trained model parameters $(A,B,\phi)$ Algorithm: 1. Initialize $A=I$, $B=\mathbf{0}$ 2. for each training epoch do 3.  for each batch of sequences do 4.   for each sequence $\{\mathbf{x}_{t}^{i}\}_{t\in\{0,\ldots,T-1\}}$ in batch do 5.    Initialize $\theta_{0}=\phi(\mathbf{x}_{0})$ 6.    for $t=1$ to $T-1$ do 7.     Update $\theta_{t}=A\theta_{t-1}+B(\mathbf{x}_{t}-\mathbf{x}_{t-1})$ 8.     Compute output $\mathbf{y}_{t}=\theta_{t}(\frac{t}{T-1})$ 9.    end for 10.    Compute sequence loss $\mathcal{L}^{i}$ 11.   end for 12.   Update parameters $A,B,\phi$ using gradient descent 13.  end for 14. end for
Figure 7: Recurrent training algorithm for WARP in its non-AR form.
Recurrent Mode
@@
The recurrent training algorithm is illustrated in its non-AR setting in Figure 7. For the AR setting trained with teacher forcing, $\mathbf{x}_{t}$ in line 7 is replaced, with probability $1-p_{\text{forcing}}$, with a sample from $\mathcal{N}(\hat{\boldsymbol{\mu}}_{t},\hat{\boldsymbol{\sigma}}_{t}^{2})$ which is taken element-wise using the classic reparameterization trick as outlined in Section 2.3. The batch of sequences (lines 4 to 11) is computed in parallel using vectorization as per the implementation details below.
Convolutional Mode
@@
Like Gu et al. [2021], WARP supports a convolutional training mode where the sequence of weights is computed efficiently using Fast-Fourier Transforms (FFTs) on modern hardware Rush and Karamcheti [2022] using Theorem 1. We use the Pythonic notation $\mathbf{u}_{0:T}\triangleq\{\mathbf{u}_{t}\}_{t=0}^{T-1}\in\mathbb{R}^{T\times D_{u}}$, and the $\star$ to denote the convolution operation. The summarized convolutional training algorithm is provided in Figure 8.
@@
Substituting this into (4), we get
@@
from which the large kernel —the sequence of columns of the Kalman controllability matrix Trélat [2005]— is extracted:
@@
In practice, however, we find the assumptions of Theorem 1 too restrictive to be applicable. Indeed, with the weight-space typically larger than the input space, i.e. $D_{\theta}\gg D_{x}$, the mapping $\mathbf{u}\mapsto B\mathbf{u}$ is not surjective. For such cases, we leverage the initial network $\phi$ to enforce additional constraints into the learning process. Theorem 2 guarantees the existence of a suitable initial input difference $\Delta\mathbf{x}_{0}$ to use as input in the convolution (8).
@@
[…]
@@
The proof is straightforward by remarking that $\theta_{0}=\phi(\mathbf{x}_{0})$. Using Equation 5, we find that
@@
Since ker $\text{ker }\phi\neq\emptyset$, $\exists v$ such that $\phi(v)=0$, and since we’ve fixed $B=\nabla\phi(\mathbf{x}_{0})$, this leads to
@@
WARP’s Convolutional Training Algorithm Input: Training sequences $\{\mathbf{x}^{i}_{t}\}_{t\in\{0,\ldots,T-1\}}^{i\in\{0,\ldots,N-1\}}$ Output: Trained model parameters $(A,B,\phi)$ Algorithm: 1. Initialize $A=I$, $B=\mathbf{0}$ 2. for each training epoch do 3.  for each batch of sequences do 4.   for each sequence $\{\mathbf{x}_{t}^{i}\}_{t\in\{0,\ldots,T-1\}}$ in batch do 5.    Initialize $\theta_{0}=\phi(\mathbf{x}_{0})$ and $\Delta\mathbf{x}_{0}$ (☞ Equations 5 and 2) 6.    Compute $\theta_{0:T}=K\star\Delta\mathbf{x}_{0:T}$ (☞ Equation 8) 7.    for $t=1$ to $T-1$ do 8.     Compute output $\mathbf{y}_{t}=\theta_{t}(\frac{t}{T-1})$ 9.    end for 10.    Compute sequence loss $\mathcal{L}^{i}$ 11.   end for 12.   Update parameters $A,B,\phi$ using gradient descent 13.  end for 14. end for
Figure 8: Convolutional training algorithm for WARP, where line 6 can be computed with (inverse) FFTs and the convolution theorem. All decoding sequence steps (lines 7-9), as well as the individual sequences (the batch from lines 4-11) are processed in parallel.
A.3 Implementation Details
@@
with hyperparameter $d_{\text{lim}}>0$. Another powerful approach which has demonstrated great success in the realm of Transformers is the dynamic tanh Zhu et al. [2025] with learnable scalars $a,b,\alpha,\beta$:
@@
with $(a,\alpha)$ initialized as the largest value encountered in the training datasets, and $(b,\beta)$ both as zero. This ensures output scaling that is consistent in shape with the classical $\tanh$ activation.
@@
In some problems like MNIST, we found it not enough to constrain the root’s output within a certain bound, as the predictions kept diverging. In such cases, mechanisms like directly clipping the weights in between time steps provided an additional form of non-linearity helpful for the model. Our weight clipping strategy differs from traditional approaches discussed in continual learning contexts Elsayed et al. [2024] as it does not consider initialization:
@@
As customary with recurrent networks training with the backpropagation through time algorithm Werbos [1990], we observed the classical problem of exploding gradients Zucchet and Orvieto [2024], which was mitigated by clipping the gradient norms within a specific bound captured by $g_{\text{lim}}=10^{-7}$.
Appendix B Datasets, Baselines & Metrics
B.1 Datasets

### v2

@@ Appendix A Related Work
The problem of sequence modelling, long dominated by Transformers [89], is experiencing a renewed focus on recurrent architectures, particularly for their efficiency and unique modelling capabilities [36; 22]. Our model, the Weight-space Adaptive Recurrent Predictor (WARP), intersects with several active research areas.
@@
The concept of leveraging the weight space of neural networks is not new; for instance, optimisers and hypernetworks inherently process weight-space features [3; 104; 38], treating them as outputs of a learning algorithm. Other works explore weight features as inputs for model analysis [88; 82] or for implicitly representing data [27]. WARP distinguishes itself by directly evolving the root network’s weights as its intermediate state, without explicitly specifying a test-time loss function to minimise. This test-time regression view is similarly observed with research on linear attention [97; 98; 92] and fast weights programming [7]. In the autoregressive forecasting setting, WARP bears striking similarities with the “delta” rule, which equally updates weights based on the difference between the prediction and the target. WARP can thus be viewed as a generalisation to broader problem settings that include classification.
@@
Linear RNNs and SSMs have re-emerged as powerful tools, largely due to their parallelisable and hardware-aware training [97; 69], with impressive performance on long sequences. Notable architectures like S4 [36] and Linear Attention [49] have massively catalysed recent advancements. While WARP builds on the efficiency of linear recurrence, its core innovation lies in its unique state parametrisation — rather than solely on the recurrent mechanism — which includes non-linearities for improved expressivity.
@@
The landscape of sequence modelling is rich with innovative designs. Hybrid models like Griffin [22] merge recurrences with attention, while Movahedi et al. [69] seek to compute dense linear RNNs from diagonal ones via fixed-point transformations. Frameworks like FACTS introduce structured memories [70]. Brain-inspired architectures [110; 7], including time-varying decoder architectures [44], seek to learn evolving relationships between model inputs and outputs. WARP contributes to this evolving field by introducing a novel mechanism — viewing the RNN hidden state as the weights and biases of a time-varying root neural network — which results in non-linear self-decoding.
@@
A central debate revolves around the true state-tracking and memory capabilities of various recurrent architectures. While some SSMs and even Transformers face theoretical limitations in solving certain tasks [65; 46], improvements like incorporating negative eigenvalues in linear RNNs aim to enhance state-tracking [33]. Other works explicitly include neural memory modules so that surprising events are more memorable [10]. The growing test-time training community [98; 92] proposes to combine recurrence with associative memories for improved sequence modelling. WARP’s use of a high-dimensional weight space for its states is a direct attempt to provide richer “infinite-dimensional” [^6] The hidden state is a function which lives in an “infinite-dimensional” space. memory capacity and more expressive temporal dynamics compared to conventional compressed state representations. This has parallels with the fast weights literature [80; 7].
@@
Effective adaptation to out-of-distribution dynamics or in continual learning settings is a significant challenge [34]. For instance, standard Neural Ordinary Differential Equations [17] struggle with distribution shifts and need retraining or fine-tuning for adaptation [55; 48]. With its gradient-free formulation, WARP facilitates test-time generalisation, a problem explored in meta-learning frameworks like Neural Context Flows [73], through differentiable closed-form solvers [12], or in-context learning [91]. WARP can be viewed as a meta-learning model given its progressive refinement of a shared initialisation $\theta_{0}$ at test-time, with strong connections to amortised inference [5; ravi2019amortised]
Koopman Operators
Our method can also be viewed as an application of Koopman operator theory to sequence-to-sequence modelling. As it is the case with nonlinear dynamical systems [56], the challenge is to identify the correct set of infinite-dimensional observable functions (the Koopman eigenfunctions) that linearise the dynamics. WARP addresses this by effectively using the neural network to learn a data-driven approximation of the Koopman operator, a technique explored in modern dynamics and machine learning [62; 66].
Appendix B Methodological Details
B.1 Motivation
@@
The main motivation behind WARP (Weight-space Adaptive Recurrent Prediction) is gradient-free adaptation to out-of-distribution (OoD) settings. Relative to OoD detection which has always been a central problem in machine learning spanning decades of research interest [19; 50], OoD adaptation is a recent but growing field rich in new and stimulating ideas [4; 73]. WARP mimics the dynamics of an idealised “smooth” gradient descent as observed through a projection of a 4898-dimensional space into a 2-dimensional PCA space in LABEL:fig:warp_vs_overfit. This offers a promising avenue for OoD adaptation with minimal cost.
(b) (a) Principal components of the weight space of WARP vs. weight trajectory fitted via the Gradient Descent strategy on a single trajectory; (b) Norm of the difference between updates as we go through the time steps for WARP, and the gradient steps for Gradient Descent.
B.2 Training Algorithms
@@
Input: Training sequences $\{\mathbf{x}^{i}_{t}\}_{t\in\{0,\ldots,T-1\}}^{i\in\{0,\ldots,N-1\}}$ Output: Trained model parameters $(A,B,\phi)$ Algorithm: 1. Initialise $A=I$, $B=\mathbf{0}$, and $\phi$ 2. for each training epoch do 3.  for each batch of sequences do 4.   for each sequence $\{\mathbf{x}_{t}^{i}\}_{t\in\{0,\ldots,T-1\}}$ in batch do 5.    Initialise $\theta_{0}^{i}=\phi(\mathbf{x}_{0}^{i})$ 6.    for $t=1$ to T1T-1 do 7.     Update $\theta_{t}^{i}=A\theta_{t-1}^{i}+B(\mathbf{x}_{t}^{i}-\mathbf{x}_{t-1}^{i})$ 8.     Compute output $\mathbf{y}_{t}^{i}=\text{MLP}_{\theta_{t}^{i}}(\tau)$ (☞ Eqs. 3, 4 and 5) 9.    end for 10.    Compute sequence loss $\mathcal{L}^{i}$ using Eq. 2 11.   end for 12.   Update parameters $A,B,\phi$ using gradient descent 13.  end for 14. end for
Algorithm 1 Recurrent training algorithm for WARP in its non-AR form.
B.2.1 Recurrent Mode
The recurrent training pipeline is illustrated in its non-AR setting in Algorithm 1, where $N$ indicates the total number of instances in the training set, indexed by $i$. The quantity $\tau$ is constructed from the components in Eqs. 3, 4 and 5. Typically, $\tau$ is formed by considering normalised time alone. However, depending on the specific use case, normalised time is concatenated with pixel coordinates (for image data), or with positional encoding using sines and cosines (e.g., for time-series analysis).
  • Normalised Time. This component consists of the normalised time step, where $T$ is the total sequence length:
    $\tau=\frac{t}{T-1}.$ (3)
  • Normalised Pixel Coordinates. For image data, spatial information is encoded using normalised pixel coordinates. Given a pixel at position $(w,h)$ in an image of total size $(W,H)$, the coordinates are:
    $\tau=\left[\frac{w}{W-1},\frac{h}{H-1}\right].$ (4)
  • Positional Encoding with Sines and Cosines. The components of this matrix $\tau=PE\in\mathbb{R}^{T\times d}$ are defined as:
    $PE_{(t,k)}=\begin{cases}\sin\left(\frac{t}{C^{2j/d}}\right)&\text{if }k=2j\\ \cos\left(\frac{t}{C^{2j/d}}\right)&\text{if }k=2j+1,\end{cases}$ (5)
    where $d$ is the encoding dimension, and $C$ is a hyperparameter that controls the frequency of the sinusoidal functions [89].
@@ B.2.1 Recurrent Mode
For the AR setting trained with teacher forcing, $\mathbf{x}_{t}^{i}$ in line 7 is replaced, with probability $1-p_{\text{forcing}}$, with a sample from $\mathcal{N}(\hat{\bm{\mu}}_{t},\hat{\bm{\sigma}}_{t}^{2})$ which is taken element-wise using the classic reparametrisation trick as outlined in Section 2.3777In the main text, the superscripts $i$ were omitted for clarity.. The batch of sequences (lines 4 to 11) is processed in parallel using vectorisation as per the implementation details below.
B.2.2 Convolutional Mode
@@
Like [36], WARP supports a convolutional training mode where the sequence of weights is computed efficiently using Fast-Fourier Transforms (FFTs) on modern hardware [79] using Theorem 1. We use the Pythonic notation $\mathbf{u}_{0:T}\triangleq\{\mathbf{u}_{t}\}_{t=0}^{T-1}\in\mathbb{R}^{T\times D_{u}}$, and the $\star$ to denote the convolution operation. The summarised convolutional training algorithm is provided in Algorithm 2.
@@
Substituting this into equation 6, we get
@@
from which the large kernel —the sequence of columns of the Kalman controllability matrix [87]— is extracted:
@@
In practice, however, we find the assumptions of Theorem 1 too restrictive to be applicable. Indeed, with the weight space typically larger than the input space, i.e. $D_{\theta}\gg D_{x}$, the mapping $\mathbf{u}\mapsto B\mathbf{u}$ is not surjective. For such cases, we leverage the initial network $\phi$ to enforce additional constraints into the learning process. Theorem 2 guarantees the existence of a suitable initial input difference $\Delta\mathbf{x}_{0}$ to use as input in the convolution equation 10.
@@
Fix $\phi$ as a locally linear operator with $B=\nabla\phi(\mathbf{x}_{0})$, and assume $\emph{ker }\phi\neq\emptyset$. There exists $v\in\mathbb{R}^{D_{x}}$ such that $\Delta\mathbf{x}_{0}=\mathbf{x}_{0}-v$ and Eq. 10 holds.
@@
The proof is straightforward by remarking that $\theta_{0}=\phi(\mathbf{x}_{0})$. Using Eq. 7, we find that
@@
Since $\text{ker }\phi\neq\emptyset$, $\exists v$ such that $\phi(v)=0$, and since we’ve fixed $B=\nabla\phi(\mathbf{x}_{0})$, this leads to
@@
Input: Training sequences $\{\mathbf{x}^{i}_{t}\}_{t\in\{0,\ldots,T-1\}}^{i\in\{0,\ldots,N-1\}}$ Output: Trained model parameters $(A,B,\phi)$ Algorithm: 1. Initialise $A=I$, $B=\mathbf{0}$, and $\phi$ 2. for each training epoch do 3.  for each batch of sequences do 4.   for each sequence $\{\mathbf{x}_{t}^{i}\}_{t\in\{0,\ldots,T-1\}}$ in batch do 5.    Initialise $\theta_{0}^{i}=\phi(\mathbf{x}_{0}^{i})$ and $\Delta\mathbf{x}_{0}$ (☞ Eqs. 7 and 2) 6.    Compute $\theta_{0:T}^{i}=K\star\Delta\mathbf{x}_{0:T}^{i}$ (☞ Eq. 10) 7.    for $t=1$ to T1T-1 do 8.     Compute output $\mathbf{y}_{t}^{i}=\text{MLP}_{\theta_{t}^{i}}(\tau)$ 9.    end for 10.    Compute sequence loss $\mathcal{L}^{i}$ using Eq. 2 11.   end for 12.   Update parameters $A,B,\phi$ using gradient descent 13.  end for 14. end for
Algorithm 2 Convolutional training algorithm for WARP, where line 6 can be computed with (inverse) FFTs and the convolution theorem. All decoding sequence steps (lines 7-9), as well as the individual sequences (the batch from lines 4-11) are processed in parallel.
B.3 Implementation Caveats
@@
with hyperparameter $d_{\text{lim}}>0$. Another powerful approach which has demonstrated great success in the realm of Transformers is the dynamic tanh [106] with learnable scalars $a,b,\alpha,\beta$:
@@
with $(a,\alpha)$ initialised as the largest value encountered in the training datasets, and $(b,\beta)$ both as zero. This ensures output scaling that is consistent in shape with the classical $\tanh$ activation.
@@
In some problems like MNIST, we found it not enough to constrain the root’s output within a certain bound, as the predictions kept diverging. In such cases, mechanisms like directly clipping the weights in between time steps provided an additional form of non-linearity helpful for the model. Our weight clipping strategy differs from traditional approaches discussed in continual learning contexts [28] as it does not consider initialisation:
@@
As customary with recurrent networks training with the backpropagation through time algorithm [95], we observed the classical problem of exploding gradients [109], which was mitigated by clipping the gradient norms within a specific bound captured by $g_{\text{lim}}=10^{-7}$.
Appendix C Datasets, Baselines & Metrics
C.1 Datasets